OpenStudy (mcba1234):

If I wanted to change the y-intercept of the graph from (0,2) to (0,5), what form would you use and what would you do to make the change?

OpenStudy (mathmale):

I'm sure there are several ways to look at this situation. First of all, I need to know whether a graph was part of this problem. The problem statement mentions changing the y-intercept of the graph. (Which graph?) If you were to use or to find the slope-intercept form of the equation of a straight line, y=mx + b, then that equation would initially look like y=mx+2. Were you to keep everything except for the y-intercept the same, and to change the y-intercept to (0,5) then the appropriate equation would look like y=mx+5.

OpenStudy (mcba1234):

There was no graph that came with the problem. My notes didn't say anything about the slope intercept form. My options are standard form, vertex form, and factored form.

OpenStudy (cathyangs):

Can you write out in symbols what those 3 forms are? I'll keep helping you if you respond (not sure if you're active atm)

OpenStudy (mcba1234):

Factored form: y=a(x-r)(x-r) Vertex form: y=a(x-h)^2+k Standard form: y=ax^2+bx+c

OpenStudy (cathyangs):

From which of these is it possible to (immediately) tell what value y would be when x is zero? If I told you that x=0, you would immediately be able to tell what value (in terms of symbols) y would be at that point?

OpenStudy (mcba1234):

Would it be factored form?

OpenStudy (cathyangs):

So in factored form, can you immediately tell what value y would be when x = 0? I personally can't tell that. I think the correct answer is your standard form, because when x =0 the terms ax^2+bx cancel to 0, and you are left with just y = c. C would be the value of y when x=0, otherwise known as the....?